OpenStudy (anonymous):

what value of a makes the equation true? A.25 B.6 C.12 D.15

OpenStudy (anonymous):

Where is the equation? :)

OpenStudy (anonymous):

\[\frac{ 24 }{ 40 }=\frac{ 9 }{ ? }\]

OpenStudy (anonymous):

?=a

OpenStudy (anonymous):

We can cross multiply.

OpenStudy (anonymous):

didnt think of that

OpenStudy (anonymous):

Multiply 40 x 9

OpenStudy (anonymous):

im horrible at this but ill give it a try

OpenStudy (anonymous):

And 24 * x

OpenStudy (anonymous):

@Tylyngaming Don't worry, I have it under control. :)

OpenStudy (anonymous):

oh,ok

OpenStudy (anonymous):

ill go help someoe then

OpenStudy (anonymous):

40*9=360 then you / bye 12 and get 15?

OpenStudy (anonymous):

Well, multiply 24 * x.

OpenStudy (anonymous):

24 srry. how you cant *x

OpenStudy (anonymous):

It's 24x

OpenStudy (anonymous):

So how you have 24x = 360. Can you tell me what we have to do now?

OpenStudy (anonymous):

We have to divide, so can you divide 360 \(\div\) 24? @sierra117